# Dance to Demise – How Massive Stars May Form Dense Circumstellar Shells Before Explosion

### v2

@@ 5.2 SN 2020ywx
SN 2020ywx is the second brightest Type IIn SN observed in X-rays after SN 2010jl, with mass loss rates inferred to be M⊙ over the last yrs before CC, from multi-epoch Swift and Chandra X-ray observations, while optical and radio data give estimates between M⊙ (R. Baer-Way et al., 2025). To compare with these observed estimates, we present two of our MESA model sequences for M⊙with and in Figure 7(a). Our M⊙, model sequence exhibits late onset of dynamical ejections with mass loss rates M⊙ following a superwind phase with lower mass loss rates of M⊙ , which resembles the optical and radio behaviour of SN 2020ywx. Whereas, our M⊙, model sequence shows a longer phase with dynamical ejections leading to mass loss rates M⊙ over the last yrs before CC, as is inferred from X-ray luminosities at multiple epochs for SN 2020ywx (R. Baer-Way et al., 2025).
The CSM density constructed using the above two model sequences is presented in Figure 7(b) along with observational estimates obtained using multi-epoch X-ray mass loss rates from R. Baer-Way et al. (2025), assuming a constant shock velocity of km and a terminal CSM wind speed of measured from optical and near-infrared spectra (refer Section 3.1 in R. Baer-Way et al. 2025). The CSM densities inferred from the observed X-ray mass loss rate estimates are in close agreement with our M⊙, model sequences, while they are underpredicted by more than an order of magnitude by the M⊙, sequences. This is primarily due to the time of onset and duration of the episodes of dynamical ejections in the two models; also the X-ray mass loss rates of R. Baer-Way et al. (2025) differ by upto an order of magnitude from the optical/radio values beyond the first epoch. Figure 7(c) presents the column densities of neutral hydrogen along the line of sight for both the models compared to the measured X-ray column densities at the four epochs of Chandra X-ray measurements at days of R. Baer-Way et al. (2025). The location of the forward shock within the CSM is found by assuming a constant shock speed of km. It is readily seen that the column densities predicted using our M⊙, model fits the observed values better than the profile obtained using the M⊙, model. The reason for this discrepancy could again arise due to asymmetry
@@ 5.3 SN 2017hcc
SN 2017hcc belongs to the category of superluminous Type IIn SNe, whose progenitor is thought to have experienced extreme eruptive mass loss in the decade before explosion, with mass loss rate of M⊙ required to power the main peak in its light-curve within first days dominated by CSM interaction post-explosion (N. Smith & J. E. Andrews, 2020). This episodic mass loss is supposed to have occurred for about yrs pre-explosion in order to have created a dense CSM around the progenitor that powered the SN luminosity through strong interaction with this dense CSM. In Figure 8(a), we present one of our MESA model sequences with M⊙, which also exhibits episodic mass loss due to dynamical ejections over the last few decades before CC, with mass loss rates reaching above M⊙ .
To construct the CSM around the progenitor of SN 2017hcc, we use X-ray observations of P. Chandra et al. (2022) who found a lower average mass loss rate of M⊙ over the last decade pre-CC, corresponding to a post shock breakout time of days, followed by a decline in the mass loss rates to M⊙ at few days, assuming a bolometric luminosity erg, a terminal wind speed of (N. Smith & J. E. Andrews, 2020) and the velocity of the forward shock to be . Furthermore, P. Chandra et al. (2022) also present radio observations from which they estimate a mass loss rate at days to be M⊙ .
Figure 8(b) shows a comparison of our model predictions with the CSM densities obtained using the observed mass loss rates from P. Chandra et al. (2022) as described above. It can be seen that our model predictions closely trace the observed density profile, even with lower average mass loss rates during both the superwind and dynamical mass ejection phases. While LBV-type eruptive mass loss over the last decade could also produce a similar density profile (B. Kumar et al., 2019), we demonstrate that dynamical ejections following a superwind phase can create a dense enough CSM as is needed to power the peak luminosity of SN 2017hcc.
In Figure 8(c), we present the neutral hydrogen column density calculated using the predicted CSM density profile of Figure 8(b), along with column density measured from the single epoch of Chandra observations of SN 2017hcc, as presented in P. Chandra et al. (2022). Our model prediction agrees with the observed column density within the uncertainties of its measured value.
@@ 5.4 SN 2005ip
SN 2005ip is the one of the best studied Type IIn SN with a long-lasting phase of strong interaction with its CSM hinting towards high mass loss rates ( M⊙ ) over the last yrs before CC (O. D. Fox et al., 2020). Eruptive mass loss from LBVs have been invoked in the literature to explain the origin of the dense shell around SN 2005ip (N. Smith et al., 2009; S. Katsuda et al., 2014). Here, we illustrate that dynamical ejections following a strong superwind phase from pulsating RSGs could also lead to enduring phases of such heavy mass loss, as is needed to reconcile with measured column densities from Swift and Chandra X-ray observations of SN 2005ip (S. Katsuda et al., 2014; O. D. Fox et al., 2020).
Figure 9(a) shows two sequences of MESA models both with M⊙, but with exhibiting such a sustained period of enhanced mass loss from dynamical ejections following a strong superwind phase with lower mass loss rates between M⊙ . The dynamical ejections occur with an average mass loss rate M⊙ over several decades before CC, which leads to M⊙of CSM material located close to the progenitor, as seen in Figure 9(b). Our model predictions closely trace the observationally inferred CSM density, computed using a mass loss rate of M⊙ (S. Katsuda et al., 2014) and a terminal wind speed of (N. Smith et al., 2009), out to distances of cm from the progenitor surface which is the location of the forward shock at days post-explosion, assuming the shock velocity and has a value of at days (O. D. Fox et al., 2020). At later times, corresponding to larger distances, the mass loss rates are estimated to be much lower at M⊙ which is also reflected in our model densities caused due to matter lost in the superwind phase preceding dynamical ejections. The model with , which has an earlier onset of dynamical ejections at yrs from CC, (N. Smith et al., 2009), shows this drop in CSM density at a location closer to the observed extent of the dense shell around the progenitor (O. D. Fox et al., 2020).
Figure 9(c) (N. Smith et al., 2009), shows the column density profile of neutral hydrogen along the line of sight obtained using the CSM densities according to 4.3.3. The calculated column densities from the model with are in close agreement with available estimates from late-time Swift/Chandra X-ray observations of S. Katsuda et al. (2014) and O. D. Fox et al. (2020), which are converted into radial coordinates from post-explosion times using the time evolution of the shock wavefront, as mentioned above.
@@ 5.5 SN 1998S
SN 1998S is considered to be the prototypical Type-IIn SN, with a mass loss rate estimate of M⊙ during the last yrs before CC and a much lower rate at earlier times (I. Shivvers et al., 2015), though J. Mauerhan & N. Smith (2012) showed substantial matter present as far out as cm from late-time spectroscopic observations yrs post-explosion. Previously, the terminal velocity of the wind and the mass loss rate of the progenitor had been estimated to be km/s and M⊙ respectively (A. Fassia et al., 2001; D. Pooley et al., 2002). Based on these studies, we use the following observed mass loss rate profile:
@@
The bottom panel of Figure 10 compares the CSM density predicted from the MESA models using Equation 5 with the values inferred from the observed mass loss rate profile given by Equation (8). The “inner” CSM ( cm) predictions closely resemble observationally estimated density profile for . Other evolutionary scenarios invoked to explain the observed light curve of SN 1998S, that requires a dense CSM nearby, include merger of a massive RSG with a compact object (S. L. Schrøder et al., 2020). However, the presence of two distinct shells as described above poses tension with this merger-driven explosion scenario, though episodic mass ejections may still occur during the common-envelope (CE) phase in such systems (M. Clayton et al., 2017).
@@ 6.1 Observational implications for the RSG problem
Our MESA evolutionary sequences with the PDSW and subsequent dynamical ejections for initial masses M⊙, presented in Section 4, undergo late-stage heavy mass loss, creating the dense CSM around the progenitor star as is required to explain observations of the Type II SNe shown in Section 5. S.-C. Yoon & M. Cantiello (2010) suggested that, considering their PDSW mass loss prescription, their exists a critical mass, M⊙ above which RSGs lose a large fraction of their H-rich envelopes prior to CC, without producing Type IIP SNe, reconciling with the so-called missing RSG problem inferred from observations of RSGs (S. J. Smartt et al., 2009; S. J. Smartt, 2015; B. Davies & E. R. Beasor, 2020; C. S. Kochanek, 2020; E. R. Beasor et al., 2025; Q. Fang et al., 2025). We performed some calculations with higher initial mass in the above range, most of which led to runaway mass loss due to shock-driven mass ejections shortly following the onset of PDSW, in the very early stages of core-carbon burning. We could not evolve these runs further due to numerical convergence issues, but such runaway events also favor the scenario in which most of the H-rich envelope lost much before CC, in agreement with the observed lack of higher mass RSGs associated with Type IIP SNe.
@@ 6.2 Pulsation-driven mass loss in binary systems
Large-amplitude pulsations can also drive enhanced mass loss in binary systems, both through PDSWs of the kind considered in this study, as well as mass transfer via pulsational Roche lobe overflow (S. Mohamed & P. Podsiadlowski, 2007), by changing the Roche lobe radius and its impact on mass transfer in the presence of pulsation-driven winds (T. Dermine et al., 2009). M. Clayton et al. (2017) also investigated the possibility of dynamical mass ejections during the inspiral of a compact object within the CE of a merging binary system, demonstrating the viability of envelope ejection following episodes of strong compression. Such scenarios may also have profound impacts on the binary evolution outcomes of more massive stars – the majority of which are thought to have experienced at least one episode of interaction with a companion star during the course of their evolution (H. Sana et al., 2012). Hence, a systematic study of the effects of both pulsation-driven mass loss and shock-driven dynamical mass ejections is warranted in the context of massive binaries to quantify the impact on pre-SN evolution as well as CSM formation.
Incorporating these mass loss mechanisms might also help to alleviate the existing tension between pre-explosion mass loss rates inferred from observations of Type Ib/c SNe (M. R. Drout et al., 2016) and those measured from stripped progenitor stars in binaries (Y. Götberg et al., 2023). Furthermore, it has been proposed late mass-transfer episodes in binaries may be stable if they are pulsed, possibly as a consequence of pulsation instabilities in the envelopes of the giant donors in such systems (P. P. Eggleton & C. A. Tout, 1989) (for detailed estimates of all probable massive binary evolutionary channels leading to various types of SNe, refer P. Podsiadlowski et al. 1992).
@@ 6.3 Effects of rotation and metallicity
The effect of rotation on the pulsational instability of RSG envelopes was examined by A. Heger et al. (1997) who found dramatic increases in ratios due to the physical effects of rotation on the evolution of massive stars, even at typical values of surface rotational velocities of which amounts to of critical rotation. This effect is primarily due to an increase in helium core masses in the evolutionary models, which translates to higher luminosities during advanced phases of evolution. This property leads to a significant increase in the growth rates of the fundamental mode of the pulsational instability, which in turn would lead to further enhancement in the strength of the associated PDSW (as per Equation 2). S.-C. Yoon & M. Cantiello (2010) concluded that rotation could further lower the effective upper mass limit () of RSGs which are expected to explode as Type IIPs, in closer agreement with latest studies on the observed value of this limit (E. R. Beasor & N. Smith, 2022; E. R. Beasor et al., 2025). This would also have consequences for the formation of Wolf-Rayet (WR) stars that are thought to be single star progenitors of Type Ib/c SNe (N. Langer & S. E. Woosley, 1996), as more rapidly rotating RSGs might already show strong pulsations during core helium burning and lose sufficient mass to evolve into WR stars (Gräfener, G. et al., 2012). It would also imply lower initial masses for WR stars than predicted from standard non-rotating evolutionary models, thereby increasing the number of single star progenitors of Type Ib/c SNe (S. E. Woosley et al., 1993).
@@
Metallicity also plays a key role in driving mass loss through line-driven winds (A. Heger et al., 2005; Mokiem, M. R. et al., 2007; L. Muijres et al., 2012), with weaker winds predicted at lower metallicities (I. Baraffe & M. F. El Eid 1991; E. R. Higgins et al. 2025; though see N. Mauron & E. Josselin 2011; K. Antoniadis et al. 2025) due to lack of metal lines needed to absorb sufficient number of photons which can impart necessary momentum to drive the wind. Due to this effect, it has been proposed that massive stars only evolve into cool RSGs above a critical metallicity (P.-S. Ou et al., 2023), below which they can advance through core helium or carbon burning while retaining most of their compact envelopes (P.-S. Ou & K.-J. Chen, 2025). I. Baraffe et al. (2001) studied the pulsational stability of very massive stars of varying metallicity and found a transition for the mechanism driving these pulsations with increasing metallicity, with the -mechanism dominating for metal-rich counterparts of primordial stars that may end their lives without losing much mass (D. Bahena & J. Klapp, 2010). This is attributed to the much longer growth timescales of nuclear-powered pulsational instabilities in metal-free stars on the main sequence, which quickly stabilize after evolving off the main sequence. However, Moriya, Takashi J. & Langer, Norbert (2015) demonstrated that pulsationally driven winds could be an efficient way for very massive stars to lose a substantial part of the envelopes in the RSG phase before explosion, even at zero metallicity. Furthermore, the prospect of dynamical mass loss following subsequent shock breakout in the outer layers of envelopes of RSGs in such pristine regimes, also deserves attention, particularly in the context of dust formation in the high-redshift universe (T. H. Greif, 2015) as observed by large-scale surveys such as the JWST and the LSST.

### v3

@@ 5.2 SN 2020ywx
SN 2020ywx is the second brightest Type IIn SN observed in X-rays after SN 2010jl, with mass loss rates inferred to be M⊙ over the last yrs before CC, from multi-epoch Swift and Chandra X-ray observations, while optical and radio data give estimates between M⊙ (Baer-Way_2025). To compare with these observed estimates, we present two of our MESA model sequences for M⊙with and in the top panel of Figure 7. Our M⊙, model sequence exhibits late onset of dynamical ejections with mass loss rates M⊙ following a superwind phase with lower mass loss rates of M⊙ , which resembles the optical and radio behaviour of SN 2020ywx. Whereas, our M⊙, model sequence shows a longer phase with dynamical ejections leading to mass loss rates M⊙ over the last yrs before CC, as is inferred from X-ray luminosities at multiple epochs for SN 2020ywx (Baer-Way_2025).
The CSM density constructed using the above two model sequences is presented in the middle panel of Figure 7 along with observational estimates obtained using multi-epoch X-ray mass loss rates from Baer-Way_2025, assuming a constant shock velocity of km and a terminal CSM wind speed of measured from optical and near-infrared spectra (refer Section 3.1 in Baer-Way_2025). The CSM densities inferred from the observed X-ray mass loss rate estimates are in close agreement with our M⊙, model sequences, while they are underpredicted by more than an order of magnitude by the M⊙, sequences. This is primarily due to the time of onset and duration of the episodes of dynamical ejections in the two models; also the X-ray mass-loss rates of Baer-Way_2025 differ by upto an order of magnitude from the optical/radio values beyond the first epoch.
The bottom panel of Figure 7 presents the column densities of neutral hydrogen along the line of sight for
both the models compared to the measured X-ray column densities at the four epochs of Chandra X-ray measurements at days of Baer-Way_2025. The location of the forward shock within the CSM is found by assuming a constant shock speed of km. It is readily seen that the column densities predicted using our M⊙, model fits the observed values better than the profile obtained using the M⊙, model. The reason for this discrepancy could again arise due to asymmetry effects or line-of-sight effects influencing the column density measurements of Baer-Way_2025, as discussed for SN 2023ixf (Chandra_2024; 2025ApJ...985...51A).
@@ 5.3 SN 2017hcc
SN 2017hcc belongs to the category of superluminous Type IIn SNe, whose progenitor is thought to have experienced extreme eruptive mass loss in the decade before explosion, with mass loss rate of M⊙ required to power the main peak in its light-curve within first days dominated by CSM interaction post-explosion (Smith&Andrews2020). This episodic mass loss is supposed to have occurred for about yrs pre-explosion in order to have created a dense CSM around the progenitor that powered the SN luminosity through strong interaction with this dense CSM. In the top panel of Figure 8, we present one of our MESA model sequences with M⊙, which also exhibits episodic mass loss due to dynamical ejections over the last few decades before CC, with mass loss rates reaching above M⊙ .
To construct the CSM around the progenitor of SN 2017hcc, we use X-ray observations of 2022MNRAS.517.4151C who found a lower average mass loss rate of M⊙ over the last decade pre-CC, corresponding to a post shock breakout time of days, followed by a decline in the mass loss rates to M⊙ at few days, assuming a bolometric luminosity erg, a terminal wind speed of (Smith&Andrews2020) and the velocity of the forward shock to be . Furthermore, 2022MNRAS.517.4151C also present radio observations from which they estimate a mass loss rate at days to be M⊙ .
The middle panel of Figure 8 shows a comparison of our model predictions with the CSM densities obtained using the observed mass loss rates from 2022MNRAS.517.4151C as described above. It can be seen that our model predictions closely trace the observed density profile, even with lower average mass loss rates during both the superwind and dynamical mass ejection phases. While LBV-type eruptive mass loss over the last decade could also produce a similar density profile (10.1093/mnras/stz1914), we demonstrate that dynamical ejections following a superwind phase can create a dense enough CSM as is needed to power the peak luminosity of SN 2017hcc.
In the bottom panel of Figure 8, we present the neutral hydrogen column density calculated using the predicted CSM density profile shown in the middle panel of Figure 8, along with column density measured from the single epoch of Chandra observations of SN 2017hcc, as presented in 2022MNRAS.517.4151C. Our model prediction agrees with the observed column density within the uncertainties of its measured value.
@@ 5.4 SN 2005ip
SN 2005ip is the one of the best studied Type IIn SN with a long-lasting phase of strong interaction with its CSM hinting towards high mass loss rates ( M⊙ ) over the last yrs before CC (Fox2020). Eruptive mass loss from LBVs have been invoked in the literature to explain the origin of the dense shell around SN 2005ip (Smith_2009; Katsuda_2014). Here, we illustrate that dynamical ejections following a strong superwind phase from pulsating RSGs could also lead to enduring phases of such heavy mass loss, as is needed to reconcile with measured column densities from Swift and Chandra X-ray observations of SN 2005ip (Katsuda_2014; Fox2020).
The top panel of Figure 9 shows two sequences of MESA models both with M⊙, but with exhibiting such a sustained period of enhanced mass loss from dynamical ejections following a strong superwind phase with lower mass loss rates between M⊙ . The dynamical ejections occur with an average mass loss rate M⊙ over several decades before CC, which leads to M⊙of CSM material located close to the progenitor, as seen in the middle panel of Figure 9. Our model predictions closely trace the observationally inferred CSM density, computed using a mass loss rate of M⊙ (Katsuda_2014) and a terminal wind speed of (Smith_2009), out to distances of cm from the progenitor surface which is the location of the forward shock at days post-explosion, assuming the shock velocity and has a value of at days (Fox2020). At later times, corresponding to larger distances, the mass loss rates are estimated to be much lower at M⊙ (Smith_2009), which is also reflected in our model densities caused due
to matter lost in the superwind phase preceding dynamical ejections. The model with , which has an earlier onset of dynamical ejections at yrs from CC, shows this drop in CSM density at a location closer to the observed extent of the dense shell around the progenitor (Fox2020).
The bottom panel of Figure 9 (Smith_2009), shows the column density profile of neutral hydrogen along the line of sight obtained using the CSM densities according to 4.3.3. The calculated column densities from the model with are in close agreement with available estimates from late-time Swift/Chandra X-ray observations of Katsuda_2014 and Fox2020, which are converted into radial coordinates from post-explosion times using the time evolution of the shock wavefront, as mentioned above.
@@ 5.5 SN 1998S
SN 1998S is considered to be the prototypical Type-IIn SN, with a mass loss rate estimate of M⊙ during the last yrs before CC and a much lower rate at earlier times (2015ApJ...806..213S), though 2012MNRAS.424.2659M showed substantial matter present as far out as cm from late-time spectroscopic observations yrs post-explosion. Previously, the terminal velocity of the wind and the mass loss rate of the progenitor had been estimated to be km/s and M⊙ respectively (2001MNRAS.325..907F; Pooley_2002). Based on these studies, we use the following observed mass loss rate profile:
@@
The bottom panel of Figure 10 compares the CSM density predicted from the MESA models using Equation 5 with the values inferred from the observed mass loss rate profile given by Equation (8). The ”inner” CSM ( cm) predictions closely resemble the observationally estimated density profile for . Other evolutionary scenarios invoked to explain the observed light curve of SN 1998S, which requires a dense CSM nearby, include merger of a massive RSG with a compact object (2020ApJ...892...13S). However, the presence of two distinct shells as described above poses tension with this merger-driven explosion scenario, though episodic mass ejections may still occur during the common-envelope (CE) phase in such systems (2017MNRAS.470.1788C).
@@ 6.1 Observational implications for the RSG problem
Our MESA evolutionary sequences with the PDSW and subsequent dynamical ejections for initial masses M⊙, presented in Section 4, undergo late-stage heavy mass loss, creating the dense CSM around the progenitor star as is required to explain observations of the Type II SNe shown in Section 5. 2010ApJ...717L..62Y suggested that, considering their PDSW mass loss prescription, their exists a critical mass, M⊙ above which RSGs lose a large fraction of their H-rich envelopes prior to CC, without producing Type IIP SNe, reconciling with the so-called missing RSG problem inferred from observations of RSGs (2009MNRAS.395.1409S; 2015PASA...32...16S; davies_2020; kochanek_2020; beasor_2025; 2025ApJ...986...39F). We performed some calculations with higher initial mass in the above range, most of which led to runaway mass loss due to shock-driven mass ejections shortly following the onset of PDSW, in the very early stages of core-carbon burning. We could not evolve these runs further due to numerical convergence issues, but such runaway events also favor the scenario in which most of the H-rich envelope lost much before CC, in agreement with the observed lack of higher mass RSGs associated with Type IIP SNe.
@@ 6.2 Pulsation-driven mass loss in binary systems
Large-amplitude pulsations can also drive enhanced mass loss in binary systems, both through PDSWs of the kind considered in this study, as well as mass transfer via pulsational Roche lobe overflow (2007ASPC..372..397M), by changing the Roche lobe radius and its impact on mass transfer in the presence of pulsation-driven winds (2009A&A...507..891D). 2017MNRAS.470.1788C also investigated the possibility of dynamical mass ejections during the inspiral of a compact object within the common envelope (CE) of a merging binary system, demonstrating the viability of envelope ejection following episodes of strong compression. Such scenarios may also have profound impacts on the binary evolution outcomes of more massive stars – the majority of which are thought to have experienced at least one episode of interaction with a companion star during the course of their evolution (2012Sci...337..444S). Hence, a systematic study of the effects of both pulsation-driven mass loss and shock-driven dynamical mass ejections is warranted in the context of massive binaries to quantify the impact on pre-SN evolution as well as CSM formation.
Incorporating these mass loss mechanisms might also help to alleviate the existing tension between pre-explosion mass loss rates inferred from observations of Type Ib/c SNe (2016ApJ...821...57D) and those measured from stripped progenitor stars in binaries (2023ApJ...959..125G). Furthermore, it has been proposed late mass-transfer episodes in binaries may be stable if they are pulsed, possibly as a consequence of pulsation instabilities in the envelopes of the giant donors in such systems (10.1007/978-94-009-2413-0_15) (for detailed estimates of all probable massive binary evolutionary channels leading to various types of SNe, refer 1992ApJ...391..246P).
@@ 6.3 Effects of rotation and metallicity
The effect of rotation on the pulsational instability of RSG envelopes was examined by 1997A&A...327..224H who found dramatic increases in ratios due to the physical effects of rotation on the evolution of massive stars, even at typical values of surface rotational velocities of which amounts to of critical rotation. This effect is primarily due to an increase in helium core masses in the evolutionary models, which translates to higher luminosities during advanced phases of evolution. This property leads to a significant increase in the growth rates of the fundamental mode of the pulsational instability, which in turn would lead to further enhancement in the strength of the associated PDSW (as per Equation 2). 2010ApJ...717L..62Y concluded that rotation could further lower the effective upper mass limit () of RSGs which are expected to explode as Type IIPs, in closer agreement with latest studies on the observed value of this limit (beasor_2022; beasor_2025). This would also have consequences for the formation of Wolf-Rayet (WR) stars that are thought to be single star progenitors of Type Ib/c SNe (1996ASPC...98..220L), as more rapidly rotating RSGs might already show strong pulsations during core helium burning and lose sufficient mass to evolve into WR stars (rsg_wr2012). It would also imply lower initial masses for WR stars than predicted from standard non-rotating evolutionary models, thereby increasing the number of single star progenitors of Type Ib/c SNe (1993ApJ...411..823W).
@@
Metallicity also plays a key role in driving mass loss through line-driven winds (Heger_2005; refId0; 2012A&A...546A..42M), with weaker winds predicted at lower metallicities (1991A&A...245..548B; 2025A&A...699A..71H; though see 2011A&A...526A.156M; 2025arXiv250305876A) due to lack of metal lines needed to absorb sufficient number of photons which can impart necessary momentum to drive the wind. Due to this effect, it has been proposed that massive stars only evolve into cool RSGs above a critical metallicity (2023ApJ...944...34O), below which they can advance through core helium or carbon burning while retaining most of their compact envelopes (2025arXiv250601753O). 2001ApJ...550..890B studied the pulsational stability of very massive stars of varying metallicity and found a transition for the mechanism driving these pulsations with increasing metallicity, with the -mechanism dominating for metal-rich counterparts of primordial stars that may end their lives without losing much mass (2010Ap&SS.327..219B). This is attributed to the much longer growth timescales of nuclear-powered pulsational instabilities in metal-free stars on the main sequence, which quickly stabilize after evolving off the main sequence. However, ml2015 demonstrated that pulsationally driven winds could be an efficient way for very massive stars to lose a substantial part of the envelopes in the RSG phase before explosion, even at zero metallicity. Furthermore, the prospect of dynamical mass loss following subsequent shock breakout in the outer layers of envelopes of RSGs in such pristine regimes, also deserves attention, particularly in the context of dust formation in the high-redshift universe (2015ComAC...2....3G) as observed by large-scale surveys such as the JWST and the LSST.